Williamson, Inc., manufactures digital voice recorders. During 2016, total costs associated with manufacturing 202,500 of the new EZ 9900 model (introduced this year) were as follows:

Raw materials $ 1,780,200
Direct labor 2,959,800
Variable manufacturing overhead 752,200
Fixed manufacturing overhead 706,900

Required:

a. Calculate the cost per recorder under both variable costing and absorption costing. (Round your answers to 2 decimal places.)

Cost per recorder
Variable costing each
Absorption costing each

b. If 20,230 of these recorders were in finished goods inventory at the end of 2016, by how much and in what direction (higher or lower) would 2016 cost of goods sold and operating income be different under variable costing than under absorption costing? (Round your intermediate calculations to 2 decimal places and final answer to the whole number.)

Total fixed cost associated with ending inventory   
Under variable costing, the cost of goods sold would be    than under absorption costing.

c-1. Express the digital voice recorder cost in a cost formula. (Round your answers to 2 decimal places.)

Total cost = + per digital voice recorder produced.

c-2. What does this formula suggest the total cost of making an additional 1,750 recorders would be?

Total additional cost
Answer #1
a) cost per recorder
variable costing
Raw materials 1780200
direct labor 2959800
variable manufacturing OH 752200
Total cost of production 5492200
Cost per recorder 5492200/202500 27.12
Absorption costing
Raw materials 1780200
direct labor 2959800
variable manufacturing OH 752200
Fixed manufacturing OH 706900
TOTAL COST 6199100
cost per recorder (6199100/202500) 30.61
b) Cost of goods sold -absorption costing (30.61*20230) = 619240.3
cost of goods sold - variable costing (27.12*20230) = 548637.6
Difference = 70602.7
under variable costing finished goods will be lower by 70602.7
c) y= a+bx
y= total cost, a = fixed cost, b= variable cost, x= units
c2) y= 0+1750*27.12
y = 47460
additional unit cost = 47460
y = 706900+(202500+1750)*27.12
y = 706900+ 5539260
y = 6246160